Calculate the acceleration of josh riding his bicycle in a straight line that speeds up from 4m/s to 6m/s in 5 seconds



1
Expert's answer
2022-02-10T08:36:03-0500

Given:

"\Delta v=v_f-v_i=\rm 2\: m/s"

"\Delta t=5\:\rm s"


The acceleration

"a=\frac{\Delta v}{\Delta t}=\frac{2\:\rm m/s}{5\:\rm s}=0.4\:\rm m/s^2"